A projectile is fired straight upward from the ground with an initial speed of 96 feet per second, then its height h after t seconds is given by the equation
h(t) = -16t2 + 96t
Find the maximum height of the projectile.

Answer by Alan3354(69443)

A projectile is fired straight upward from the ground with an initial speed of 96 feet per second, then its height h after t seconds is given by the equation
h(t) = -16t^2 + 96t
Max is the vertex of the parabola, at t = -b/2a
t = -96/-32 = 3 seconds
--------
h(3) = -16*9 + 96*3 = 144 feet